\begin{thebibliography}{99}

\bibitem{as}
\textsc{M. Abramowitz, I. Stegun, eds.}:
\textit{Handbook of mathematical functions with formulas,
 graphs and mathematical tables.}  National Bureau of Standards, Dover, New York, 1965.


\bibitem{aq}
\textsc{H. Alzer, S.-L. Qiu }:
\textit{Monotonicity theorems and inequalities for the generalized complete elliptic integrals,} J. Comput. Appl. Math., \textbf{172} 2004, No. 2, pp:289-312.


\bibitem{ar}
\textsc{H. Alzer, K. Richards }:
\textit{A note on a function involving complete elliptic integrals: Monotonicity, convexity, inequalities,} Anal. Math., \textbf{41} 2015, pp: 133-139.


\bibitem{aqv}
\textsc{G.D. Anderson, S.-L. Qiu, and M.K. Vamanamurthy}:
\textit{Elliptic integrals inequalities, with applications,}
Constr. Approx., \textbf{14}, 1998, pp:195-207.


\bibitem{avvb}
\textsc{G.D. Anderson, M.K. Vamanamurthy, and M. Vuorinen}:
\textit{Conformal invariants, inequalities and quasiconformal
maps.}  J. Wiley, 1997, 505 pp.




\bibitem{av}
\textsc{G.D. Anderson, and M.K. Vamanamurthy}:
\textit{ Some properties of quasiconformal distortion functions,}
New Zealand J. Math., \textbf{24}, 1995, pp:1-16.



\bibitem{aar}
\textsc{G.E. Andrews, R. Askey and R. Roy}:
\textit{Special functions.}  Cambridge University Press, Cambridge, 1999.

\bibitem{adv}
\textsc{G.D. Anderson, P. Duren, and M.K. Vamanamurthy}:
\textit{ An inequality for elliptic integrals,}
J. Math. Anal. Appl., \textbf{182}, 1994, pp:257-259.


\bibitem{bbk}
\textsc{\'{A}. Baricz, B.A. Bhayo, R. Kl\'en}: 
\textit{Convexity properties of generalized trigonometric and hyperbolic functions}, Aequat. Math. 89 (2015), 473--484.


\bibitem{bv}
\textsc{B.A. Bhayo, M. Vuorinen}: 
\textit{On generalized trigonometric functions  with two parameters}, 
J. Approx. Theory 164 (2012), 1415--1426.



\bibitem{bhayoyin}
\textsc{B.A. Bhayo, and L. Yin}:
\textit{On generalized $(p,q)$-elliptic integrals},
arXiv:1507.00031 [math.CA]



\bibitem{bb}
\textsc{J.M. Borwein and P.B. Borwein}:
\textit{Pi and the AGM, A study in analytic
number theory and computational complexity}, Reprint of the 1987 original. Canadian Mathematical Society Series of Monographs and Advanced Texts, 4. A Wiley-Interscience Publication. John Wiley \& Sons,
Inc., New York, 1998


\bibitem{brent}
\textsc{R.P. Brent}:
\textit{Fast multiple-precision evaluation of elementary functions,} J. Assoc. Comput. Math. 23 (1976), no. 2, 242--251.

\bibitem{be}
\textsc{P.J. Bushell, D.E. Edmunds}: 
\textit{Remarks on generalised trigonometric functions}, Rocky Mountain J. Math. 42 (2012), 13--52.


\bibitem{cg}
\textsc{F. Carlson and J.L. Gustafson}:
\textit{ Asymptotic approximations for symmetric elliptic integrals,}
SIAM J. Math. Anal., \textbf{25}, 1994, 288--303.


\bibitem{ref2}
\textsc{Y.-M. Chu and M.-K. Wang}:
\textit{Optimal Lehmer mean bounds for the
Toader mean}, Results Math., 2012, 61(3-4), 223--229.


\bibitem{ref3}
\textsc{Y.-M. Chu, M.-K.Wang and S.-L. Qiu}:
\textit{Optimal combinations bounds
of root-square and arithmetic means for Toader mean}, Proc. Indian Acad.
Sci. Math. Sci., 2012, 122(1), 41--51.


\bibitem{ref4}
\textsc{Y.-M. Chu, Y.-F. Qiu and M.-K. Wang}:
\textit{H\"older mean inequalities
for the complete elliptic integrals}, Integral Transforms Spec. Funct., 2012,
23(7), 521--527.


\bibitem{ref5}
\textsc{Y.-M. Chu, M.-K. Wang, S.-L. Qiu and Y.-P. Jiang}:
\textit{Bounds for complete
elliptic integrals of the second kind with applications}, Comput. Math.
Appl., 2012, 63(7), 1177--1184.



\bibitem{dm}
\textsc{P. Dr\'abek and R. Man\'asevich}:
\textit{On the closed solution to some $p-$Laplacian nonhomogeneous eigenvalue problems}.
Diff. and Int. Eqns. 12 (1999), 723-740.



\bibitem{egl}
\textsc{D.E. Edmunds, P. Gurka, J. Lang}:
\textit{ Properties of generalized trigonometric functions}, J. Approx. Theory 164 (2012), 47--56.




\bibitem{ref1}
\textsc{T.-R. Huang, S.-Y. Tan, X.-Y. Ma and Y.-M. Chu}: 
\textit{Monotonicity properties
and bounds for the complete p-elliptic integrals}, J. Inequal. Appl.,
2018, Article 239, 11 page



\bibitem{HTZ}
\textsc{T.-R. Huang,S.-Y. Tan, X.-H. Zhang}: 
\textit{Monotonicity, convexity, and inequalities for the generalized elliptic integrals}. J. Inequal. Appl. 2017, Article ID 278 (2017).



%\bibitem{hvv}
%\textsc{V. Heikkala, M. K. Vamanamurthy and M. Vuorinen}:
%\textit{Generalized elliptic integrals,} Comput. Methods Funct. Theory, \textbf{9}, 2009, No. 1, pp:75-109.
%
%\bibitem{l}
%\textsc{P. Lindqvist}:
%\textit{Some remarkable sine and cosine functions},
%Ricerche di Matematica Vol. XLIV (1995), 269-290.




\bibitem{pbm}
\textsc{A.P. Prudnikov, Y.A. Brychkov and O.I. Marichev}:
\textit{Integrals and series,} V 3, Gordon and Breach Science Publishers(Amsterdam-New York, 1990).


\bibitem{ref9}
\textsc{W.-M. Qian and Y.-M. Chu}:
\textit{Sharp bounds for a special quasi-arithmetic
mean in terms of arithmetic and geometric means with two parameters}, J.
Inequal. Appl. 2017, Article 274, 10 pages.

\bibitem{salamin}
\textsc{E. Salamin}:
\textit{Computation of $\pi$ using arithmetic-geometric means},
Math. Comp. 30 (1976), no. 135, 565--570.


\bibitem{wzc}
\textsc{G.-D. Wang, X.-H. Zhang and Y.-M. Chu}:
\textit{Inequalities for the generalized elliptic integrals and modular equations,} J. Math. Anal.
Appl., \textbf{331} 2007, No. 2, pp:1275-1283.

\bibitem{t1}
\textsc{S. Takeuchi}:
\textit{Generalized Jacobian elliptic functions and their application to bifurcation problems associated with $p$-Laplacian,}
J. Math. Anal. Appl. \textbf{385}, 2012, pp:24-35.

\bibitem{t2}
\textsc{T. Kamiya and S. Takeuchi}:
\textit{Complete $(p,q)$-elliptic integrals with application to a family of means,}
J. Class. Anal. 10 (2017), no. 1, 15--25.

\bibitem{t3}
\textsc{S. Takeuchi}:
\textit{The complete $p$-elliptic integrals and a computation formula of $\pi_p$ for $p=4$},
Ramanujan J. 46 (2018), no. 2, 309--321.

\bibitem{t4}
\textsc{S. Takeuchi}:
\textit{A new form of the generalized complete elliptic integrals},
 Kodai Math. J. 39 (2016), no. 1, 202--226.

\bibitem{wg}
\textsc{Zh.-X. Wang and D.-R. Guo}:
 \textit{Introduction to Special Function,} Peking University Press,
Beijing, 2004.



\bibitem{ref6}
\textsc{M.-K. Wang, S.-L. Qiu and Y.-M. Chu}:
\textit{Infinite series formula for H\"ubner upper bound function with applications to Hersch-Puger distortion
function}, Math. Inequal. Appl., 2018, 21(3), 629--648.


\bibitem{ref7}
\textsc{M.-K. Wang, Y.-M. Li and Y.-M. Chu}:
\textit{Inequalities and infinite product
formula for Ramanujan generalized modular equation function}, Ramanujan J., 2018, 46(1), 189--200.


\bibitem{ref8}
\textsc{M.-K. Wang and Y.-M. Chu}:
\textit{Landen inequalities for a class of hypergeometric
functions with applications}, Math. Inequal. Appl., 2018, 21(2),
52--537.


\bibitem{ref10}
\textsc{M.-K. Wang and Y.-M. Chu}:
\textit{Refinements of transformation inequalities
for zero-balanced hypergeometric functions}, Acta Math. Sci., 2017,
37B(3), 607--622.


\bibitem{ref11}
\textsc{M.-K. Wang, Y.-M. Chu, Y.-F. Qiu and S.-L. Qiu}:
\textit{An optimal
power mean inequality for the complete elliptic integrals}, Appl. Math. Lett.,
2011, 24(6), 887--890.
%%%%%%%%%%%%%%%

\bibitem{ref12}
\textsc{Zh.-H. Yang, W.-M. Qian and Y.-M. Chu}:
\textit{Monotonicity properties
and bounds involving the complete elliptic integrals of the first kind}, Math.
Inequal. Appl., 2018, 21(4), 1185--1199.

\bibitem{ref13}
\textsc{Zh.-H. Yang, W.-M. Qian, Y.-M. Chu and W. Zhang}:
\textit{On approximating
the arithmetic-geometric mean and complete elliptic integral of the first
kind}, J. Math. Anal. Appl., 2018, 462(2), 1714--1726.

\bibitem{ref14}
\textsc{Zh.-H. Yang and Y.-M. Chu}:
\textit{A monotonicity property involving the
generalized elliptic integral of the first kind}, Math. Inequal. Appl., 2017,
20(3), 729--735.

\bibitem{ref15}
\textsc{Zh.-H. Yang, W.-M. Qian, Y.-M. Chu and W. Zhang}:
\textit{Monotonicity
rule for the quotient of two functions and its application}, J. Inequal. Appl.,
2017, Article 106, 13 pages.

\end{thebibliography}

\documentclass[12pt,a4paper,leqno,verbatim]{amsart}
\newcounter{minutes}
\divide\time by 60
\newcounter{hours}
\multiply\time by 60 \addtocounter{minutes}{-\time}

\usepackage{amssymb}
\usepackage{hyperref}
\usepackage{graphicx}
\usepackage{subfigure}
\date{}
\newfont{\cyrilic}{wncyr10 scaled 1000}

\title[{Generalized complete elliptic integrals}]
{On a function involving generalized complete \\
$(p,q)$- elliptic integrals}

\author[B. A. Bhayo]{Barkat Ali Bhayo}
\address{Department of mathematics, Sukkur IBA University, Sindh, Pakistan}
\address{Faculty of Engineering and natural sciences, Sabanci University, 34956 Tuzla/Istanbul, Turkey}
\email{barkat.bhayo@iba-suk.edu.pk,\quad barkatbhayo@sabanciuniv.edu}

\author[L. Yin]{Li Yin}
\address{Department of Mathematics, Binzhou University, Binzhou City, Shandong Province, 256603, China}
\email{yinli\_79@163.com}

\newcommand{\comment}[1]{}

\swapnumbers
\theoremstyle{plain}

\newtheorem{theorem}[equation]{Theorem}
\newtheorem{lemma}[equation]{Lemma}

\newtheorem{remark}[equation]{Remark}

\newtheorem{proposition}[equation]{Proposition}

\numberwithin{equation}{section}

\begin{document}

\font\fFt=eusm10 
\font\fFa=eusm7  
\font\fFp=eusm5  
\def\K{\mathchoice
{\hbox{\,\fFt K}}
{\hbox{\,\fFt K}}
{\hbox{\,\fFa K}}
{\hbox{\,\fFp K}}}
\def\E{\mathchoice
{\hbox{\,\fFt E}}
{\hbox{\,\fFt E}}
{\hbox{\,\fFa E}}
{\hbox{\,\fFp E}}}

\allowdisplaybreaks

\begin{abstract}
Motivated by the work of Alzer and Richards \cite{ar}, here authors study the monotonicity and convexity properties of the function
$$\Delta _{p,q} (r) = \frac{{E_{p,q}(r)  - \left( {r'} \right)^p K_{p,q}(r) }}{{r^p }} - \frac{{E'_{p,q}(r)  - r^p K'_{p,q}(r) }}{{\left( {r'} \right)^p }},$$
where $K_{p,q}$ and $E_{p,q}$ denote the complete $(p,q)$- elliptic integrals of the first and second kind, respectively.
\end{abstract}

\vspace{.5cm}

\maketitle
{\bf 2010 Mathematics Subject Classification}: 33C99, 33B99

{\bf Keywords and phrases}: inequalities, $(p,q)-$ elliptic integrals, hypergeometric function, monotonicity, convexity.


\def\thefootnote{}
\footnotetext{ \texttt{\tiny File:~\jobname .tex,
          printed: \number\year-\number\month-\number\day,
          \thehours.\ifnum\theminutes<10{0}\fi\theminutes}
} \makeatletter\def\thefootnote{\@arabic\c@footnote}\makeatother

\section{introduction}

For $0<r<1$ and $r'=\sqrt{1-r^2}$, the Legendre's complete elliptic integrals of the first and the second kind are defined by

$$\left\{\begin{array}{lll} \K=\K(r)=\displaystyle\int_0^{\pi/2}\frac{dt}{\sqrt{1-r^2\sin(t)^2}}=
\displaystyle\int_0^1{\frac{dt}{\sqrt{(1-t^2)(1-r^2t^2)}}},\\

\bigskip

 \E=\E(r)=\displaystyle\int_0^{\pi/2}\sqrt{1-r^2
\sin(t)^2}dt=\displaystyle\int_0^1\sqrt{\frac{1-r^2t^2}{1-t^2}}dt,\\
                      \K(0)=\displaystyle\frac{\pi}{2}=
											\E(0),\, \K(1)=\infty,
											\,\E(1)=0,\\
											\K'=\K'(r)=\K(r')\quad {\rm and }\quad \E'=\E'(r)=\E(r'),
\end{array}\right.$$
respectively. These integrals have played very crucial role in many branches of mathematics, for example, they helps us to find the length of curves and to express the solution of differential equations. The elliptic integrals of the first and the second kind have been extensive interest of the research for several authors, and many results have been established about these integrals in the literature, e.g., 
see \cite{ref2,ref3,ref4,ref5,ref1,ref9,ref11}.
For the monotonicity, convexity properties, asymptotic approximations, functional inequalities of these integrals and their relations with elementary functions, we refer the reader to see e.g., \cite{avvb,ref6,ref7,ref8,ref10,ref11,ref12,ref13,ref14,ref15} and the references therein.

The Gaussian hypergeometric function is defined by
\begin{equation}\label{hypform}
F(a,b;c;z)=~_{2}F_1(a,b;c;z)=\sum_{n\geqslant 0}\frac{(a)_n(b)_n}{(c)_n}\frac{z^n}{n!}, \quad |z|<1.
\end{equation}
Here $(a)_0=1$ for $a\neq 0$, and $(a)_n$ is the Pochhammer symbol $(a)_n=a(a+1)\cdots(a+n-1)$, for $n\in \mathbf{N}$.
The Gaussian hypergeometric function can be represented in the integral form as follows,
\begin{equation}\label{hyp}
F(a,b;c;z)=\frac{\Gamma(c)}{\Gamma(b)(c-b)}\int_0^1{t^{b-1}(1-t)^{c-b-1}(1-zt)^{-a}dt},
\end{equation}
${\rm Re}(c)>{\rm Re}(a)>0, |\arg(1-z)|<\pi$, see \cite{as}.

For $|s|<1/2$ and $0\leq|r|<1$, the complete elliptic integrals of the first and the 
second kind were slightly generalized by Borwein and Borwein \cite{bb} as follows:
\begin{equation}\label{start}{\rm K}_s(r)=F\left(\frac{1}{2}-s,\frac{1}{2}+s;1;r^2\right),\quad
{\rm E}_s(r)=F\left(-\frac{1}{2}-s,\frac{1}{2}+s;1;r^2\right).
\end{equation}
Note that ${\rm K}_0(r)=\K(r)\quad {\rm and}\quad {\rm E}_0(r)=\E(r).$

In order to define the generalized complete $(p,q)$-elliptic integrals of the first and the second kind, we need to define the generalized sine function.

The eigenfunction $\sin_{p,q}$ of the so-called one-dimensional $(p,q)$-Laplacian problem
\cite{dm}
$$-\Delta_p u=-\left(|u'|^{p-2}u'\right)'
=\lambda|u|^{q-2}u,\,u(0)=u(1)=0,\ \ \ p,q>1,$$
is known as the generalized sine function with two parameters $p,q>1$ in the literature 
(see, \cite{bbk,bv,be,egl,t1,t2,t3,t4}), and defined as
the inverse function of
$${\rm arcsin}_{p,q}(x)=\int^x_0(1-t^p)^{-\frac{1}{q}}dt,\quad 0<x<1.$$
Also the generalized $\pi$ is defined as
$$\pi_{p,q}=2{\rm arcsin}_{p,q}(1)=\frac{2}{q}B\left(1-\frac{1}{p},\frac{1}{q}\right),$$
which is the generalized version of the celebrated formula of $\pi$ proved by Salamin \cite{salamin} and Brent \cite{brent} in 1976.
Here $B(.,.)$ denotes the classical beta function.

For all $p,q\in (1,\infty),\,r\in (0,1)$ and $r' = \left( {1 - r^p } \right)^{1/p}$, the generalized complete $(p,q)$-elliptic integrals of the first and the second kind are defined by
\begin{equation}\label{buffer}
\begin{cases}
K_{p,q} (r) = \int_0^{\pi _{p,q} /2} {\left( {1 - r^q \sin_{p,q} ^{q} t} \right)^{1/p-1}} dt,\quad
K'_{p,q}  = K'_{p,q} (r) = K_{p,q} (r'),\\
E_{p,q} (r) = \int_0^{\pi _{p,q} /2} {\left( {1 - r^q \sin_{p,q} ^{q} t} \right)^{1/p} } dt,
\quad E'_{p,q}  = E'_{p,q} (r) = E_{p,q} (r'),
\end{cases}
\end{equation}
respectively.
Applying the integral representation formula \eqref{hyp}, the generalized complete $(p,q)$-elliptic integrals can be expressed in terms of
hypergeometric functions as follows:
\begin{equation}\label{(1.1)}
K_{p,q} (r) = \frac{{\pi _{p,q} }}{2}F\left( {\frac{1}{q},1 - \frac{1}{p};1- \frac{1}{p}+\frac{1}{q};r^p } \right),
\end{equation}
and
\begin{equation}\label{(1.2)}
E_{p,q} (r) = \frac{{\pi _{p,q} }}{2}F\left( {\frac{1}{q}, - \frac{1}{p};1- \frac{1}{p}+\frac{1}{q};r^p } \right),
\end{equation}
see, e.g. \cite{bhayoyin}.
For $p=q$, we write $K_{p,p}=K_p$. Note that $K_2=\K$ and $E_2=\E$. It is worth to mention that
Takeuchi \cite{t2} proved that
$${\rm K}_s(r)=\frac{\pi}{\pi_p}K_p(r^{2/p})\quad {\rm and} \quad
{\rm E}_s(r)=\frac{\pi}{\pi_p}E_p(r^{2/p}),$$
for $|s|<1/2$ and $p=2/(2s+1)$.


In 1998, Anderson, Qiu and Vamanamurthy \cite[Theorem 1.14]{aqv} studied the monotonicity and convexity property of the function
$$f(r)=\frac{\E-r'^2\K}{r^2}\cdot\frac{r'^2}{\E'-r^2\K'}$$
by giving the following theorem.

\begin{theorem}\label{thm1.5} The function $f(r)$ is increasing and convex from $(0,1)$ onto
$(\pi/4,4/\pi)$. In particular,
$$\frac{\pi}{4}<f(r)<\frac{\pi}{4}+\left(\frac{4}{\pi}-\frac{\pi}{4}\right)r$$
for $r\in (0,1)$. These two inequalities are sharp as $r\to 0$, while the second inequality
is also sharp as $r \to 1$.
\end{theorem}

Theorem \ref{thm1.5} has been extended to the case of the functions  \eqref{start} in \cite[Theorem 3.1]{HTZ} and to the case of 
functions $K_{p,p},\, E_{p,p}$ \eqref{buffer} in \cite[Theorem 3.1]{ref1}.

Recently, Alzer and Richards \cite{ar} studied the properties of the additive counterpart
$$\Delta (r) = \frac{{\E - \left( {1 - r^2 } \right)\K}}{{r^2 }} - \frac{{\E' - r^2 \K'}}
{{\left( {1 - r^2 } \right)}}$$
of the above result, and proved the following theorem.

\begin{proposition}\label{alri}
The function $\Delta (r)$ is strictly increasing and strictly convex from $(0,1)$ onto $(\pi/4-1,1-\pi/4)$. Moreover, for all $r\in(0,1)$, one has
\begin{equation}\label{(1.3)}
\frac{\pi }{4} - 1 + \alpha r < \Delta (r) < \frac{\pi }{4} - 1 + \beta r,
\end{equation}
with best possible constants $\alpha=0$ and $\beta=2-\frac{\pi}{2}=0.42920\ldots.$
\end{proposition}

It is natural to extend the result of Alzer and Richards in terms of generalized complete $(p,q)-$ elliptic integrals of the first and second kind. We generalize their function $\Delta$ by
$$\Delta _{p,q} (r) = \frac{{E_{p,q}  - \left( {r'} \right)^p K_{p,q} }}{{r^p }} - \frac{{E'_{p,q}  - r^p K'_{p,q} }}{{\left( {r'} \right)^p }},
$$
and state the following theorem.

\begin{theorem}\label{thm1.1}
For $p,q>1$, the function $
\Delta_{p,q}$ is strictly increasing and strictly convex from $(0,1)$ onto $
\left( {\frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}} - 1,1 - \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}}} \right)$,
if the following conditions hold:
\begin{enumerate}
\item $2 + \frac{1}{p} + \frac{1}{{p^2 }} \le \frac{5}{p} + \frac{1}{q} < 3 + \frac{1}{{p^2 }},$ \\
\item $\varepsilon(p,q)>0,$
where $$\varepsilon(p,q)=
20 - \frac{{42}}{p} + \frac{6}{q} + \frac{{21}}{{p^2 }} - \frac{2}{{q^2 }} - \frac{{20}}{{pq}} + \frac{9}{{p^2 q}} - \frac{3}{{p^3 }} - \frac{1}{{p^3 q}}.$$
\end{enumerate}

Moreover, for all $r\in(0,1)$, we have
\begin{equation}\label{(1.4)}
{\frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}} - 1}
+ \alpha_1r < \Delta_{p,q} (r) <
{\frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}} - 1}
+ \beta_1 r,
\end{equation}
with best possible constants $\alpha_1=0$ and $\beta_1=
{2 - \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}}}.$
\end{theorem}

\begin{theorem}\label{thm1.2}
For all $r,s\in(0,1)$ and $p,q>1$ satisfying the conditions (1) and (2) given in the above theorem, then we have 
\begin{equation}\label{(1.5)}
{\frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}} - 1}
 < \Delta _{p,q} (rs) - \Delta _{p,q} (r) - \Delta _{p,q} (s) <
{1 - \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}}}.
\end{equation}
\end{theorem}

\section{Lemmas}
In this section we give few lemmas which will be used in the proof of the theorems. Moreover, we will use same method for proving our theorems as it is applied in \cite{ar}.

\begin{lemma}\label{lem2.1} Write
$$
H_{a,b} (r) = \frac{\pi _{1/b,1/a} }{2r^{1/b} }\left[ F\left( {a, - b;1+a-b;r^{1/b} } \right) -
\left( 1-r^{1/b} \right) F\left( {a,1 - b;1+a-b;r^{1/b} } \right) \right].$$

For $a,b,r\in(0,1)$, we have
\begin{equation}\label{(2.2)}
H_{a,b} (r) = \frac{(1 - b)\pi _{1/b,1/a} }{2(1+a-b)}F\left( a,1 - b;2+a-b;r^{1/b}  \right).
\end{equation}
\end{lemma}

\begin{proof}
By using the formula \eqref{hypform}, we obtain
$$
\begin{array}{l}
 H_{a,b} (r) =\displaystyle\frac{\pi _{1/b,1/a} }{2r^{1/b} }\left[ {\sum\limits_{n = 0}^\infty  {\frac{{(a)_n ( - b)_n }}{{(1+a-b)_n{n!}  }}r^{n/b} }  - \left( {1 - r^{1/b} } \right)\sum\limits_{n = 0}^\infty  {\frac{{(a)_n (1 - b)_n }}{{(1+a-b)_n{n!} }}r^{n/b} } } \right] \\
  =\displaystyle\frac{{\pi _{1/b,1/a} }}{{2r^{1/b} }}\left[ {\sum\limits_{n = 0}^\infty  {\left( {\frac{{(a)_n ( - b)_n }}
	{{(1+a-b)_n{n!} }} - \frac{{(a)_n (1 - b)_n }}{{(1+a-b)_n{n!} }}} \right)r^{n/b} }  + \sum\limits_{n = 0}^\infty  {\frac{{(a)_n (1 - b)_n }}{{(1+a-b)_n{n!} }}r^{(n + 1)/b} } } \right] \\
  =\displaystyle\frac{{\pi _{1/b,1/a} }}{2} \sum\limits_{n = 0}^\infty  {\left( {\frac{{(a)_{n + 1} ( - b)_{n + 1} }}{{(1+a-b)_{n+1}{(n+1)!} }} - \frac{{(a)_{n + 1} (1 - b)_{n + 1} }}{{(1+a-b)_{n+1}{(n+1)!} }}} \right)r^{n/b} }     \\
	\qquad +\displaystyle\frac{{\pi _{1/b,1/a} }}{2}\sum\limits_{n = 0}^\infty  {\frac{{(a)_n (1 - b)_n }}{{(1+a-b)_{n}{n!} }}r^{n/b} }\\
  =\displaystyle\frac{{\pi _{1/b,1/a}} }{2}\left[ {\sum\limits_{n = 0}^\infty  {\frac{{(a)_n (1 - b)_n }}{{(1+a-b)_{n+1}{(n+1)!} }}\cdot \xi_{a,n} \cdot r^{n/b} } } \right] \\
  =\displaystyle\frac{{(1 - b)\pi _{1/b,1/a} }}{2}\sum\limits_{n = 0}^\infty  {\frac{{(a)_n (1 - b)_n }}{{n!(1+a-b)_{n+1} }}r^{n/b} }  \\
  =\displaystyle\frac{{(1 - b)\pi _{1/b,1/a} }}{2(1+a-b)}F\left( {a,1 - b;2+a-b;r^{1/b} } \right), \\
 \end{array}$$
where
$$\xi_{a,n} = (n + a)( -b) - (a + n)(1 - b + n)
	+ (n + 1)(n+1+a-b)=(1-b)(n+1).$$
\end{proof}


\begin{lemma}\label{lem2.4}
For $p,q>1$ and $r\in(0,1)$, we have
\begin{equation}\label{(2.7)}
H_{1/q,1/p} (0) = \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{2\left(1 + \frac{1}{q} - \frac{1}{p}\right)},\qquad H_{1/q,1/p} (1) = 1.
\end{equation}
\end{lemma}
\begin{proof}
By definition, it is easy to see that $$
H_{1/q,1/p} (0) = \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{2\left(1 + \frac{1}{q} - \frac{1}{p}\right)}F\left( {\frac{1}{q},1 - \frac{1}{p};2+\frac{1}{q} - \frac{1}{p};0} \right) =  \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{2\left(1 + \frac{1}{q} - \frac{1}{p}\right)}.
$$
Again, by using the following identity
\begin{equation}\label{(2.8)}
F\left( {\alpha ,\beta ;\gamma ;1} \right) = \frac{{\Gamma (\gamma )\Gamma (\gamma  - \alpha  - \beta )}}{{\Gamma \left( {\gamma  - \alpha } \right)\Gamma \left( {\gamma  - \beta } \right)}},
\end{equation}
(see, \cite[p. 153]{wg}), we get
\begin{align*}
 H_{1/q,1/p} (1) = &\frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{2\left(1 + \frac{1}{q} - \frac{1}{p}\right)}F\left( {\frac{1}{q},1 - \frac{1}{p};2+\frac{1}{q} - \frac{1}{p};1} \right) \\
  = &\frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{2\left(1 + \frac{1}{q} - \frac{1}{p}\right)}
  \frac{{\Gamma \left(2+ \frac{1}{q} - \frac{1}{p}\right)\Gamma (1)}}{{\Gamma \left( {2 - \frac{1}{p}} \right)\Gamma \left( {1 + \frac{1}{q}} \right)}} \\
  =&\frac{{\pi _{p,q} }}{2}\frac{{\Gamma \left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}}{{\Gamma \left( {1 - \frac{1}{p}} \right)\Gamma \left( {1 + \frac{1}{q}} \right)}} \\
  =& 1.\\
 \end{align*}
\end{proof}

\begin{lemma}\label{lem2.3}
For $p,q>1$ and $r\in(0,1)$, we have
\begin{equation}\label{(2.5)}
\begin{array}{l}
 \left( {2 - \frac{1}{p}} \right)F\left( {1 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};1 - r^p } \right)\\
  =\left(3 + \frac{1}{q} - \frac{1}{p}\right)F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};1 - r^p } \right) \\
  - \left( {1 + \frac{1}{q}} \right)F\left( {1 + \frac{1}{q},2 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};1 - r^p } \right). \\
 \end{array}
\end{equation}
\end{lemma}
\begin{proof}
Utilizing the contiguous relation (See \cite[Equation 26]{pbm})
$$
(\sigma  - \rho )F\left( {\alpha ,\rho ;\sigma  + 1;z} \right) = \sigma F\left( {\alpha ,\rho ;\sigma ;z} \right) - \rho F\left( {\alpha ,\rho  + 1;\sigma  + 1;z} \right),
$$
and letting $
\sigma  = 3 + \frac{1}{q} - \frac{1}{p},\alpha  = 1 + \frac{1}{q},\rho  = 2 - \frac{1}{p},z = 1 - r^p
$, we get the identity  ~\eqref{(2.5)}.
\end{proof}

\section{Proof of the main result}
\vspace{.5cm}
\noindent{\bf Proof of Theorem \ref{thm1.1}.}
Using formulas ~\eqref{(1.1)},  ~\eqref{(1.2)} and letting $a=1/q,\,b=1/p$ in \eqref{(2.2)}, we have
\begin{align*}
 \Delta _{p,q} (r) =& H_{1/q,1/p} (r) - H_{1/q,1/p} (r') \\
  = &\frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{2\left(1 + \frac{1}{q} - \frac{1}{p}\right)}\left[ {F\left( {\frac{1}{q},1 - \frac{1}{p};2 + \frac{1}{q} - \frac{1}{p};r^p } \right) - F\left( {\frac{1}{q},1 - \frac{1}{p};2 + \frac{1}{q} - \frac{1}{p};1 - r^p } \right)} \right]. \\
\end{align*}
Applying the following derivative formula$$
\frac{d}{{dr}}F\left( {a,b;c;r} \right) = \frac{{ab}}{c}F\left( {a + 1,b + 1;c + 1;r} \right),
$$
we obtain
$$\Delta '_{p,q} (r) =
 \eta _{p,q} r^{p - 1} \left[ {F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};r^p } \right) + F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right)} \right],$$
and
$$
\begin{array}{l}
 \frac{1}{{\eta _{p,q} r^{p - 2} }}\Delta ''_{p,q} (r) =  \\
 (p - 1)\left[ {F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};r^p } \right) + F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right)} \right] \\
  + \frac{{p\left( {1 + \frac{1}{q}} \right)\left( {2 - \frac{1}{p}} \right)}}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)}}r^{p}\left[ {F\left( {2 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};r^p } \right) - F\left( {2 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right)} \right], \\
 \end{array}
$$ where
$$\eta _{p,q}  = \frac{{\frac{p}{q}\left( {1 - \frac{1}{p}} \right)^2 \pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)\left( {2 + \frac{1}{q} - \frac{1}{p}} \right)}}.$$
By utilizing the following identity
$$(1 - z)F\left( {a + 1,b + 1;a + b + 1;z} \right) = F\left( {a,b;a + b + 1;z} \right)$$
(see \cite{ar}) and letting $a=1+\frac{1}{q},b=2-\frac{1}{p},z=1-r^p$, we get
$$
\begin{array}{l}
 \frac{1}{{\eta _{p,q} r^{p - 2} }}\Delta ''_{p,q} (r) =  \\
 (p - 1)\left[ {F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};r^p } \right) - F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right)} \right] \\
  + \frac{{p\left( {1 + \frac{1}{q}} \right)\left( {2 - \frac{1}{p}} \right)}}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)}}r^p F\left( {2 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};r^p } \right) \\
  - \frac{{p\left( {1 + \frac{1}{q}} \right)\left( {2 - \frac{1}{p}} \right)}}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)}}F\left( {1 + \frac{1}{q},2 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right). \\
 \end{array}
$$
Now by Lemma~\ref{lem2.3} we have
$$
\begin{array}{l}
 \frac{1}{{\eta _{p,q} r^{p - 2} }}\Delta ''_{p,q} (r) =  \\
 (p - 1)F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};r^p } \right) +(p-1)F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right) \\
  + \frac{{p\left( {1 + \frac{1}{q}} \right)\left( {2 - \frac{1}{p}} \right)}}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)}}r^p F\left( {2 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};r^p } \right) \\
  - p\left( {2 - \frac{1}{p}} \right)F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right) \\
  + \frac{{p\left( {2 - \frac{1}{p}} \right)^2 }}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)}}F\left( {1 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right) \\
  = (p - 1)F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};r^p } \right) \\
  + \frac{{p\left( {1 + \frac{1}{q}} \right)\left( {2 - \frac{1}{p}} \right)}}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)}}r^p F\left( {2 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};r^p } \right) \\
  - \left[ {p\left( {2 - \frac{1}{p}} \right) - (p - 1)} \right]F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right) \\
  + \frac{{p\left( {2 - \frac{1}{p}} \right)^2 }}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)}}F\left( {1 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right). \\
 \end{array}
$$
Considering the condition (1) and (2) we conclude that
$$
\begin{array}{l}
 \frac{1}{{\eta _{p,q} r^{p - 2} }}\Delta ''_{p,q} (r) \\
  \ge (p - 1) - pF\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right) \\
  + \frac{{p\left( {2 - \frac{1}{p}} \right)^2 }}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)}}F\left( {1 + \frac{1}{q},3 - \frac{1}{p};4 + \frac{1}{q} - \frac{1}{p};\left( {r'} \right)^p } \right) \\
	
  = (p - 1)\left[ {1 + \sum\limits_{n = 0}^\infty  {\frac{{\left( {1 + \frac{1}{q}} \right)_n \left( {2 - \frac{1}{p}} \right)_n}}
	{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)_{n+1}}}\left( {n - \frac{{\frac{1}{q} + \frac{3}{p} - 1 - \frac{1}{{p^2 }}}}{{1 - \frac{1}{p}}}} \right)\frac{{\left( {r'} \right)^{pn} }}{{n!}}} } \right] \\

  > (p - 1)\left[ {1 + \sum\limits_{n = 0}^1 {\frac{{\left( {1 + \frac{1}{q}} \right)_n \left( {2 - \frac{1}{p}} \right)_n}}{{\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)_{n+1}}}\left( {n - \frac{{\frac{1}{q} + \frac{3}{p} - 1 - \frac{1}{{p^2 }}}}{{1 - \frac{1}{p}}}} \right)\frac{{\left( {r'} \right)^{pn} }}{{n!}}} } \right] \\
	
  = (p - 1)\left[ {\frac{{\varepsilon (p,q)}}{{\left( {1 - \frac{1}{p}} \right)\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)\left( {4 + \frac{1}{q} - \frac{1}{p}} \right)}} + \frac{{\left( {1 + \frac{1}{q}} \right)\left( {2 - \frac{1}{p}} \right)\left( {\frac{4}{p} + \frac{1}{q} - 2 - \frac{1}{{p^2 }}} \right)r^p }}{{\left( {1 - \frac{1}{p}} \right)\left( {3 + \frac{1}{q} - \frac{1}{p}} \right)\left( {4 + \frac{1}{q} - \frac{1}{p}} \right)}}} \right] \\
  > 0.
 \end{array}
$$
Now we conclude that $\Delta '_{p,q} (r) $ is strictly increasing on $(0,1)$, because $\Delta ''_{p,q} (r)>0 $. Hence, $\Delta '_{p,q} (r)> \Delta '_{p,q}(0).$
By L'H\^{o}spital rule, we get
\begin{align*}
 \Delta _{p,q} ^\prime  (0) =& \mathop {\lim }\limits_{r \to 0^ +  } \frac{{\Delta _{p,q} (r) - \Delta _{p,q} (0)}}{{r - 0}} \\
  =& \mathop {\lim }\limits_{r \to 0^ +  }\left[ \frac{{H_{1/q,1/p} (r) - H_{1/q,1/p} (0)}}{{r }} - \frac{{H_{1/q,1/p} (r') - H_{1/q,1/p} (1)}}{{r }}\right] \\
  =& H'_{1/q,1/p} (0) - \mathop {\lim }\limits_{x \to 1^ +  } \frac{{H_{1/q,1/p} (x) - 1}}{{\left( {1 - x^p } \right)^{1/p} }} \\
  = &\mathop {\lim }\limits_{x \to 1^ +  } H'_{1/q,1/p} (x)\frac{{\left( {1 - x^p } \right)^{1 - 1/p} }}{{x^{p - 1} }} \\
  = &0,
\end{align*}
where we apply $$
\begin{array}{l}
 H'_{1/q,1/p} (0)  \\
 =\displaystyle\frac{{\frac{p}{q}\left( {1 - \frac{1}{p}} \right)^2 r^{p - 1}\pi_{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)\left( {2 + \frac{1}{q} - \frac{1}{p}} \right)}}F\left( {1 + \frac{1}{q},2 - \frac{1}{p};3 + \frac{1}{q} - \frac{1}{p};r^p } \right)\left| {r = 0}. \right. \\
 \end{array}$$
This implies that $\Delta_{p,q} (r) $ is strictly increasing on $(0,1)$.
Define $$
M_{p,q} (r) = \frac{{\Delta _{p,q} (r) - \Delta _{p,q} (0)}}{{r - 0}}.
$$
Since $\Delta_{p,q} (r) $ is strictly convex on $(0,1)$, it follows that $M_{p,q} (r)$ is strictly increasing on $(0,1)$. This leads to
\begin{equation}\label{(2.9)}
M_{p,q} (0)<M_{p,q} (r)<M_{p,q} (1).
\end{equation}
Using Lemma~\ref{lem2.4}, we have $M_{p,q} (0)=0$ and
$$M_{p,q} (1) = \Delta _{p,q} (1) - \Delta _{p,q} (0) = 2H_{1/q,1/p} (1) - 2H_{1/q,1/p} (0) =
{2 - \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}}}.
$$
So, the formula ~\eqref{(2.9)} implies the double inequalities in ~\eqref{(1.4)}.
This completes the proof.$\hfill\square$

\vspace{.4cm}
\noindent{\bf Proof of Theorem \ref{thm1.2}.}
Let $$
\lambda _{p,q} (r,s) = \Delta _{p,q} (rs) - \Delta _{p,q} (r) - \Delta _{p,q} (s).
$$
Simple computation yields $$
\frac{\partial }{{\partial r}}\lambda _{p,q} (r,s) = s\Delta '_{p,q} (rs) - \Delta _{p,q} ^\prime  (r)
$$ and $$
\frac{{\partial ^2 }}{{\partial r\partial s}}\lambda _{p,q} (r,s) = \Delta '_{p,q} (rs) + rs\Delta _{p,q} ^{\prime \prime } (rs).
$$
Considering the results of Theorem \ref{thm1.1}, we have $\frac{{\partial ^2 }}{{\partial r\partial s}}\lambda _{p,q} (r,s)>0$. So, the function
$\frac{\partial }{{\partial r}}\lambda _{p,q} (r,s)$ is strictly increasing with respect to $s$. Thus, $$
\frac{\partial }{{\partial r}}\lambda _{p,q} (r,s) < \frac{\partial }{{\partial r}}\lambda _{p,q} (r,s)\left| {_{s = 1} } \right. = 0.
$$
It follows that $
r \mapsto \lambda _{p,q} (r,s)
$ is strictly decreasing which lead to $$
 - \Delta _{p,q} (1) = \lambda _{p,q} (1,s) < \lambda _{p,q} (r,s) < \lambda _{p,q} (0,s) =-\Delta _{p,q} (s)<  - \Delta _{p,q} (0).
$$ where we apply
$$
\Delta _{p,q} (0) = H_{1/q,1/p} (0) - H_{1/q,1/p} (1) = \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}} - 1$$
and
$$
\Delta _{p,q} (1) = H_{1/q,1/p} (1) - H_{1/q,1/p} (0) = 1 - \frac{{\left( {1 - \frac{1}{p}} \right)\pi _{p,q} }}{{2\left( {1 + \frac{1}{q} - \frac{1}{p}} \right)}}.$$
This completes the proof.$\hfill\square$

\begin{remark}\rm
When $p=q=2$, then it is easy to observe that Theorem \ref{thm1.2} coincides with Proposition \ref{alri}.
\end{remark}

\end{document}